OpenStudy (erinkb99):

CHECK MY ANSWER!!!! Use the mid-point rule with n = 4 to approximate the area of the region bounded by y = x^3 and y = x.

OpenStudy (erinkb99):

x^3=x x= -1, 0, 1 y=0, -1, 1 deltax= (b-a)/n deltax= (1-(-1)/4 deltax= 2/4 = 0.5 [-1, 0] midpoint= 1/2 (-1+0) midpoint= 1/2(-1) = -0.5 [0, 1] midpoint= 1/2 (1+0) midpoint= 1/2(1) = 0.5 area[-1,1] = 0.5 (0.5+(-0.5)) area[-1,1] = 0.5 (0) = 0

OpenStudy (erinkb99):

@Evoker

OpenStudy (evoker):

I'm looking, trying to follow your work, looks like you need f(-1),f(-.5),f(0),f(.5), and f(1)

OpenStudy (evoker):

I assume for midpoint rule you want the sum of (delta x)(f(mid upper)-f(mid lower))

OpenStudy (erinkb99):

Yea

OpenStudy (evoker):

Quickly looking up the method are we using f(-1), f(-.5) or f(-.75) for the first region.

OpenStudy (evoker):

Ah midpoints so yes the first needs to be f(-.75)

OpenStudy (evoker):

so -.75 for one and -.421875 for the other which yields a difference of .328125 and multiplied time .5 yields an area for the first region of .1640625

OpenStudy (erinkb99):

yes, right so its f(-0.75), f(-0.5), f(-0.25) then f(0.25), f(0.50), f(0.75)

OpenStudy (evoker):

well four regions so f(-.75),f(-.25),f(.25),f(.75)

OpenStudy (erinkb99):

ok

OpenStudy (evoker):

Should be able to repeat what I did for the first and then add all the region together.

OpenStudy (erinkb99):

Wont it still be zero

OpenStudy (erinkb99):

[-1, 0] f(-0.75) + f(-0.25) [0, 1] f(0.75) + f(0.25) area[-1,1] = f(-0.75) + f(-0.25) + f(0.75) + f(0.25) area[-1,1] = -0.421875 + -0.0625 + 0.421875 + 0.0625 = 0

OpenStudy (evoker):

Are you doing negative area because if not you want to make the first two positive

OpenStudy (evoker):

Geometrically they all should have positive results

OpenStudy (erinkb99):

True, I'm not quite sure what they are looking for

OpenStudy (evoker):

I think you should make them all positive and then add them up or in essence twice the sum of the last two.

OpenStudy (erinkb99):

ok so positive area: area[-1,1] = f(-0.75) + f(-0.25) + f(0.75) + f(0.25) area[-1,1] = 0.421875 + 0.0625 + 0.421875 + 0.0625 = 0.96875

OpenStudy (evoker):

Yes that would be my answer

OpenStudy (erinkb99):

ok I have more questions stick around!!!!

OpenStudy (evoker):

well don't forget the .5 base for each

OpenStudy (erinkb99):

oh yea

OpenStudy (erinkb99):

0.96875 (0.5)

OpenStudy (evoker):

hmm by the way I got, .328125 for the first region before the .5 element

OpenStudy (erinkb99):

how?

OpenStudy (evoker):

Ok once again for the first region you need to take the difference of f(-.75) for the two functions.

OpenStudy (evoker):

for y=x this gives a result of -.75 while y=x^3 yields -.428175

OpenStudy (evoker):

The difference is .328125

OpenStudy (erinkb99):

Oops right

OpenStudy (evoker):

So repeating for the second region f(-.25) gives -.25 and -.015625 yielding a difference of .234375

OpenStudy (evoker):

Third region f(.25) gives .25 and .015625 yielding the same difference of .234375

OpenStudy (evoker):

And fourth will be the same as first

OpenStudy (erinkb99):

ok one minute

OpenStudy (erinkb99):

positive area: area[-1,1] = deltax (f(-0.75) + f(-0.25) + f(0.75) + f(0.25)) area[-1,1] = 0.5 ((0.75-0.421875) + (0.25-0.0625) + (0.75-0.421875) + (0.25-0.0625)) area[-1,1] = 0.5 ((0.328125) + (0.1875) + (0.328125) + (0.1875)) = 0.515626

OpenStudy (erinkb99):

@Evoker

OpenStudy (evoker):

I had instead of .1875 a result of .234375 for f(.25) and f(-.25)

OpenStudy (evoker):

.015625 for the cubed term not .0625

OpenStudy (erinkb99):

yea, I squared it

OpenStudy (erinkb99):

positive area: area[-1,1] = deltax (f(-0.75) + f(-0.25) + f(0.75) + f(0.25)) area[-1,1] = 0.5 ((0.75-0.421875) + (0.25-.015625) + (0.75-0.421875) + (0.25-.015625)) area[-1,1] = 0.5 ((0.328125) + (0.234375 ) + (0.328125) + (0.234375 )) = 0.5625

OpenStudy (erinkb99):

@Evoker

OpenStudy (evoker):

Looks good now

OpenStudy (erinkb99):

Great. Next question!

OpenStudy (evoker):

Ok